# Quantification of the Self-Excited Emotion Dynamics in Online Interactions

### v1

@@ Results
In summary, the intensity $\lambda^{e}(t)$ from Equation (1) is parametrized by a set of $142|\mathcal{E}|+2=$ 14 parameters: the exogeneous rate $\mu^{e}_{0}$, the decay time $\gamma^{e}$, as well as $|\mathcal{E}|$ parameters for $\nu^{e,f}$ and $\alpha^{e,f}$, respectively. We estimate these parameters by maximizing their log-likelihood functions (see Methods). Since each emotion has its own intensity $\lambda^{e}$, our model has a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters. However, the estimation of the different $\lambda^{e}$ are decoupled from each other, so that we can estimate each set of $12$ parameters separately. The fitted parameters are visualized in Figure 2 and discussed below.
Quantifying Exogenous and Endogenous Influences
Figure 3: (a) For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time. Plot (a) shows the distribution of these ratios across videos. We notice that emotions are predominantly triggered endogenously. (b) Same as plot (a) but for the ratio of spontaneous (video influenced) intensity to total exogenous intensity.
Videos provide a shared experience among viewers, giving rise to amplified emotion intensities during live broadcasting events [30]. However, it is unclear whether the emotional experience and amplification stem from the live videos themselves or the interactions among peers [30]. The Hawkes framework (1) allows us to answer this question by comparing exogenously and endogenously triggered events. The exogenous (exo) component, $\mu^{e}(t)=\mu^{e}_{0}+\mu_{1}^{e}(t)$, represents influences from the video content and elsewhere. The endogenous (endo) component, $\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\phi^{e,f}\left(t-t_{j}^{f}\right)$, captures the influence from prior live chat messages and hence stems from interactions with peers. At any given time, we can determine the fraction of exo-influence as $R^{e}_{\text{exo}}(t)=\left.\mu^{e}(t)\right/\lambda^{e}(t)$ and the ratio of endo influence as $R^{e}_{\text{endo}}(t)=1-R^{e}_{\text{exo}}(t)$ where all quantities are calculated with the estimated parameters (Figure 2). Subsequently, we calculate the average exo- and endo-ratios, $\langle R^{e}_{\text{exo}}\rangle$ and $\langle R^{e}_{\text{endo}}\rangle$, of a given video by averaging across all times events take place. Figure 3 (left) shows the distribution of these quantities over all videos in our sample. We observe that, despite the constant exogenous influence from video feeds, endogenous influence dominates the emotion dynamics across emotion types. This demonstrates that emotional expressions in live discussions are disproportionally driven by social interactions with active participants rather than passive video content consumption. In particular, we see that the emotion of joy has the highest ratio of endogenous influence, being 80% driven by prior participants. In contrast, and in accordance to common sense, the emotion of surprise is the most responsive to exogenous stimuli where roughly 50% of the total intensity is triggered exogenously.
Similarly, we can calculate the ratio of purely exogenous spontaneous emotions relative to video-induced emotions, that is $R^{e}_{0}=\mu^{e}_{0}/\left(\mu^{e}_{0}+\mu^{e}_{1}(t)\right)$ and its complement $R^{e}_{1}=1-R^{e}_{0}.$ Figure 3 (right) shows the distribution of $\langle R^{e}_{0}\rangle$ versus $\langle R^{e}_{1}\rangle$ over the set of all videos. The rates of chats induced by videos is typically one-fourth the rate of spontaneous chats. We note, however, that there is significant heterogeneity across video categories (SI Appendix).
In summary, these results suggest that participants use YouTube live videos more as a platform for expressing their opinions and emotions than for engaging with the content itself.
Emotion Contagion and Interaction Dynamics
We now interpret the estimated parameters from Equation (1) visualized in Figure 2.
Above, we have seen that the majority of the live chat messages are triggered endogenously. An alternative way to confirm this high level of endogeneity is directly via the estimated branching ratio matrix $\alpha^{e,f}$. The branching ratio matrix quantifies the interaction between different event types within the process. The branching ratio matrix is a square matrix where each element represents the expected number of type $e$ emotional events directly triggered by a type $f$ emotional event. The diagonal elements of the branching ratio matrix represent the self-excitation within the same emotion. The off-Diagonal elements capture cross-excitation, where an emotional event of type $f$ increases the likelihood of an emotional event of type $e$. The sum $\sum_{e}\alpha^{e,f}$ represents the expected number of offspring events generated by a single parent event of type $f$. If this sum exceeds $1$ for any event type, it implies a supercritical process for that type, leading to potentially explosive behavior (more and more events). Another important quantity is the spectral radius of the branching ratio matrix, defined as the largest absolute value of its eigenvalues. It gives an overall measure of the system’s tendency toward self-excitation and cross-excitation. If the spectral radius is smaller than $1$, the system is subcritical, meaning that the process is stable, and the number of events will not grow indefinitely. The influence of past events decays over time. This is the relevant regime found in Figure 2 with a spectral radius equal to $0.8$. Examining the entries of the branching ratio matrix shown in Figure 3 (left), we find that online emotion dynamics are driven 3-5 times more by social interactions than by exogenous influences. We thus henceforth focus on the endogenous parameters $\alpha^{e,f}$ and $\gamma^{e}$.
Zooming in on the entries of the $\alpha^{e,f}$ matrix reveals that each of the 6 emotions displays distinct temporal dynamics (Figure 2, left). One can readily notice that joy has significantly higher levels of self-excitation rate than other emotions as shown by the comparison between the different diagonal entry $\alpha^{e,e}$ representing self-excitation. Quantitatively, positivity is up to 3 times more contagious than even the high-arousal negative emotions of anger and disgust. The fact that surprise is the least contagious of all the emotions serves as a useful consistency check. Similarly, fear shows relatively low levels of excitation.
Generally, self-excitation dominates over cross-excitations that are quantified by the off-diagonal values $\alpha^{e,f}$ representing the influence of emotion $f$ on emotion $e$. A notable exception is the moral emotions of anger and disgust which are mutually reinforcing in almost equal magnitudes. It is also noteworthy that all of the emotions, but particularly disgust and sadness, trigger joy. We will come back to these observations in the Discussion section below.
Recall that the decay times $\gamma^{e}$ characterize the direct time scale over which past emotions trigger future ones of type $e$. Figure 3 shows that users have a significantly longer direct memory of negative emotions, in particular of the high-arousal emotions of anger and disgust. Coupled with the fact that emotions are triggered primarily through self-excitation, this suggests that negative emotions impose a more lasting direct influence on future interactions and the emotional state of users.
Discussion
Across 92,412 live chat messages on YouTube, we find that, despite stimulative inputs from the video, online emotions are driven more by endogenous user interactions by a factor of 3 to 5. The dynamics of emotions are primarily driven by the incentives for peer communication and are further amplified through social interactions. Our results show that endogenous user interactions are the main contributors to online emotion expressions, which has been found to be consistent predictors for user engagement on video platforms [43, 44]. This provides empirical evidence that emotions are inherently communicative, being both evoked by social interactions and expressed through communication [2]. We find that live videos, while serving as the primary medium for content delivery, have only a limited direct impact on user discussions, which largely arise spontaneously. However, live videos play a crucial role in creating a shared experience that fosters empathy and enhances social interactions [3].
By quantifying the excitation and temporal decay of different emotions in online interactions, we show that positivity is more contagious whereas negative emotions create a more enduring impact on users’ memory. This reconciles mixed findings on online emotion propagation by identifying different pathways for positive and negative emotions to spread across online platforms. We further quantify an intrinsic joyful bias in users’ heightened propensity for spontaneous expressions of joy. Our results suggest that positivity bias [20, 21, 22] and negativity bias [23, 24, 25] coexist on different dimensions of social platform interactions. Users are positively biased in the spontaneous expression of emotions, and yet negatively biased in the perception of emotions.
While self-excitation dominates, anger and disgust cross-excite one another in comparable magnitudes in online interactions. It may be, however, that this result is partially explained by the frequent co-labeling of anger and disgust in the same text (see SI Appendix). Either way, the strong self- and cross-excitation of these high arousal negative emotions can explain the frequent occurrence of large-scale moral outrage in online conversation [45], and hence contributes to increased divisiveness and polarization [46, 47].
We further observe interaction effects across positive and negative emotions. In particular, both disgust and sadness trigger joy in online interactions. Such patterns can be observed in trolling and anti-social behaviors online. Individuals are less restrained online and trolling behaviors are prevalent on social platforms [48, 49]. Prior trolling activities can further trigger trolling behavior in future interactions [50].
A notable limitation of our analysis is that it is only text-based. Future studies incorporating multi-modal inputs will allow us to understand different emotion dynamics across communication channels. We also do not distinguish human participants from bots and automated accounts in YouTube live chats. Such a distinction would allow for a direct generalization of our methodology. Motivated by the growing capability of LLMs to simulate human emotions and infiltrate online environments, future studies are needed to better understand the emotional exchange between humans and modern chatbots.
Conclusions
Using 92,412 YouTube live chat messages across 397 videos, we found that interactions among users are 3-5 times more emotionally contagious than passive video content. In particular, positivity is more contagious whereas negativity is more memorable. Such understanding of online emotion dynamics has practical implications for the design of social platforms. First, user engagement is more effectively facilitated through interactive features than passive content. Second, promoting joyful content can effectively spread positivity, while the promotion of negative emotions can have lasting and alarming consequences, such as provoking moral outrage. Moreover, with the increasing ability of large language models (LLMs) to simulate human emotions [51], our results highlight the significant potential threat posed by internet bots in manipulating collective emotions through user interactions, which could potentially disrupt social order.
Methods
Collecting YouTube Live Chats
We examine emotion dynamics in YouTube live videos with replay functionality across 27 topics (see SI Appendix for the complete list). We first compile the list of keywords and collect completed live video IDs for each keyword using the keyword search function via YouTube API in Python. 111 https://developers.google.com/YouTube/v3 Subsequently, we retrieve the video transcript and live chat messages for each video respectively. We use the ChatDownloader package 222 https://github.com/xenova/chat-downloader to retrieve the available timestamped live chat content, and the YouTube-transcript-API 333 https://pypi.org/project/YouTube-transcript-api/ to retrieve the video transcript. We then only retain videos with a median time difference between two subsequent live chat messages between 1 second and 5 minutes. We exclude videos where over 70% of live chat messages are non-English. For the remaining videos, we remove live chat messages that are not in English. For each video, we also remove any live chat messages that occur before the first or after the last transcript arrival time. This leaves us with a total of 673,551 live chat messages across 1,957 videos.
Labeling Emotions in Text
Original joy anticipation optimism love trust sadness pessimism anger disgust fear surprise
Mapped joy joy joy joy joy sadness sadness anger disgust fear surprise
Table 1: We fine-tune a transformer model to detect the presence of 6 basic emotions in text. Originally, our training data contained 11 emotion labels, we mapped the 11 emotions to 6 basic emotions prior to training according to this Table.
[…]
For each video and emotion, we calculate the average number of live chat messages per minute. To avoid outlier or otherwise irregular live chats, we only retain videos for which this average lies within the $20^{\text{th}}$- and $80^{\text{th}}$-quantile across all videos. This gives us a final sample of 92,412 live chat messages across 397 videos, which amounts to a total of 780 hours. The median duration of a video is 107 minutes and the median number of live chat messages per video is 204. Across all videos, the median number of live chat messages per minute is 1.63 for joy, 0.17 for disgust, 0.16 for anger, 0.11 for sadness, 0.03 for fear, and 0.03 for surprise. We refer to the SI Appendix for additional statistics.
@@ Methods
As elaborated above, we model the arrival of emotions as a bi-hexavariate Hawkes process given by expression (1). For emotion $e\in\mathcal{E}$, within a given live chat session, we observe $N^{e}$ events, with the $j^{\text{th}}$ event taking place at time $t_{j}^{e}$. We define $t_{0}^{e}\equiv 0$, $t_{N^{e}+1}^{e}\equiv T$ where $T$ is the video duration. This way, the observation period over the entire process is $\{t_{i}^{e}\mid i=0,\ldots,N^{e}+1\}$. Similarly, we denote by $\tau^{e}_{i}$ the time at which the $i$-th subscript of emotion $e$ appears in the video. We measure time in units of minutes.
Prior studies have shown that assuming the exogenous influence $\mu^{e}(t)$ to be constant may lead to false attribution to endogenous effects [55, 40]. The YouTube live videos provide us with the unique opportunity to observe a time-varying influence of the video content on user discussions. We use the video transcript to capture the time-varying emotional content of the video. For instance, it is reasonable to assume that a particularly sad scene in the video induces sad emotions in the live chat. Further, one can expect cross-influence, e.g. that a sad scene in the video induces angry emotions in the live chat (Figure 1). In particular, this is to be distinguished from a sad scene in the video triggering a sad message in the live chat, which in turn triggers either a sad or angry message. To capture such cross-influence from the videos, we parameterize the exogenous intensity of emotion $e$ as
@@
where $\mu^{e}_{0}$ is a time-invariant, unobserved baseline intensity of spontaneous expression of emotion $e$ in the absence of influence from video or prior live chat messages. By contrast, we denote by $S^{f}(t)$ the time-varying intensity of emotion $f$ in the video, and $\nu^{e,f}$ the cross-influence from emotion $f$ in the video to emotion $e$ in the live chat. The term $\nu^{e,f}\leavevmode\nobreak\ S^{f}(t)$ is thus the time-varying intensity at which emotions of type $f$ in the video trigger emotions of type $e$ in the live chat.
The intensity $S^{f}(t)$ is given by $S^{f}(t)=\sum_{\tau^{f}_{j}<t}s^{f}_{\tau_{j}}(t)$ where $\left\{\tau^{f}_{j}\right\}$ enumerates all times at which a subtitle of emotion $f$ first appears in the video. Here, $s^{f}_{\tau_{j}}(t)$ captures the time-varying influence of a video subtitle of emotion $f$ appearing in the video at time $\tau_{j}$. Clearly, $s^{f}_{\tau_{j}}(t)=0$ for $t<\tau_{j}$. In order to incorporate the video impact, we need to make reasonable assumptions on the shape of the temporal dependence of the influence of emotions in the video on emotions in the live chat. It is natural to assume that it takes a few seconds for the audience to process the transcript and react. We also expect the influence of the video to peak after a rapid increase in intensity and then fade out with time. Empirical evidence demonstrates that information retention in humans decays with a fat tail, i.e. slower than an exponential [56]. To approximately capture the rapid initial increase and subsequent slow decline, we use the log-normal function to parametrize the memory of a subscript in the video:
@@
To fix $\mu$ and $\sigma$, we assume that the intensity of each transcript peaks $2$ seconds after appearance, and that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance. In other words, we assume that the maximum of the log-normal function lies at $2$ seconds, $\exp(\mu-\sigma^{2})=2$, and that the median is equal to $10$ seconds, $\exp(\mu)=10$. Numerically solving for $\mu$ and $\sigma$ yields $2.3$ and $1.3$ respectively in units of minutes. To ensure that our results are not strongly dependent on this choice of $\mu$ and $\sigma$, we have checked that our results remain qualitatively similar for different values and functional shapes (see SI Appendix). Alternatively, these parameters could be directly fitted from the data, a task which we leave for future research.
In summary, the shape $S^{f}(t)$ is a sum of log-normal functions, with local peaks at $2$ seconds after the appearance of a new subtitle of emotion $f$. An example of $s^{\text{sad}}$ is shown in Figure 1(d). We further stress that our framework is different from the typical multivariate Hawkes process in that expression (2) contains cross-influence terms from different exogenous sources. This is because we condense a bi-hexavariate Hawkes model containing two systems of events into one multivariate Hawkes self-excited conditional point process. This specification is therefore necessary to distinguish video-based emotion events and chat-based emotion events.
Fitting the Parameters of the Hawkes Process
@@
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}^{e}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}(s)}-\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\phi^{e,f}(s)}.$ (4b)
We follow the common assumption that $\phi$ is exponentially decaying and write $\phi^{e,f}(t)=\alpha^{e,f}\frac{1}{\gamma^{e}}e^{-\frac{1}{\gamma^{e}}t}$ with decay time $\gamma^{e}$, such that (4b) can be expanded into
$\begin{split}\log L^{e}\left(\mu^{e}_{0},\nu^{e,f},\alpha^{e,f},\gamma^{e}\right)=\sum_{i=1}^{N^{e}}\log\left(\mu^{e}_{0}+\sum_{f\in\mathcal{E}}\nu^{e,f}S^{f}(t_{i})+\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\alpha^{e,f}\frac{1}{\gamma^{e}}e^{-\frac{1}{\gamma^{e}}(t_{i}-t_{j}^{f})}\right)-\\ \quad\mu^{e}_{0}T-\sum_{f\in\mathcal{E}}\nu^{e,f}M^{f}_{1}+\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\alpha^{e,f}\left(e^{-\frac{1}{\gamma^{e}}(T-t_{j}^{f})}-1\right),\end{split}$ (5)
@@
We estimate values for $\mu^{e}_{0}$, $\nu^{e,f}$, $\alpha^{e,f}$, and $\gamma^{e}$ by maximizing Equation (5) via Quasi-Newton optimization. We use the “L-BFGS-B” optimization algorithm from the family of quasi-Newton methods to handle bound constraints. The variable bound constraints are specified as follows: $0\leq\alpha\leq 50$, $10^{-6}\leq\nu\leq 10$, $0.1\leq\gamma\leq 20$, $0\leq\mu\leq 50$.

### v2

@@ Methods
In summary, the intensity $\lambda^{e}(t)$ from Equation (1) is parametrized by a set of $142|\mathcal{E}|+2=$ 14 parameters: the exogeneous rate $\mu^{e}_{0}$, the decay time $\gamma^{e}$, as well as $|\mathcal{E}|$ parameters for $\nu^{e,f}$ and $\alpha^{e,f}$, respectively. We estimate these parameters by maximizing their log-likelihood functions (see below). Since each emotion has its own intensity $\lambda^{e}$, our model has a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters. However, the estimations of the different $\lambda^{e}$ are decoupled from each other, so that we can estimate each set of $12$ parameters separately. The fitted parameters are visualized in Figure 2 and discussed in the Results section.
@@
We model the arrival of emotions as a multivariate Hawkes process given by expression (1). For emotion $e\in\mathcal{E}$, within a given live chat session, we observe $N^{e}$ events, with the $j^{\text{th}}$ event taking place at time $t_{j}^{e}$. We define $t_{0}^{e}\equiv 0$, $t_{N^{e}+1}^{e}\equiv T$ where $T$ is the duration of the video. In this way, the observation period for the entire process is $\{t_{i}^{e}\mid i=0,\ldots,N^{e}+1\}$. Similarly, we denote by $\tau^{e}_{i}$ the time at which the $i$-th subscript of emotion $e$ appears in the video. We measure time in units of minutes. In the rest of this subsection, we describe the process of parameterizing the time-varying exogenous video influence $S^{f}(t)$. In essence, $S^{f}(t)$ denotes the level of emotion $f$ exhibited in the video at a particular time $t$ (refer to Figure 1(d) for an illustration). Readers who are mainly interested in the qualitative results of our study may skip this and the next subsection and continue with the Results Section.
Previous studies have shown that assuming the exogenous influence $\mu^{e}(t)$ to be constant can lead to false attribution to endogenous effects [66, 60] (see SI Appendix for additional confirmation). The YouTube live videos provide us with the unique opportunity to observe a time-varying influence of video content on user discussions. We use the video transcript to capture the time-varying emotional content of the video. For example, it is reasonable to assume that a particularly sad scene in the video induces sad emotions in the live chat. In addition, cross-influence can be expected; for example, a sad scene in the video induces angry emotions in the live chat (Figure 1). In particular, this is to be distinguished from a sad scene in the video triggering a sad message in the live chat, which in turn triggers either a sad or angry message. To capture such cross-influence from the videos, we parameterize the exogenous intensity of emotion $e$ as
@@
where $\mu^{e}_{0}$ is a time-invariant, unobserved baseline intensity of spontaneous expression of emotion $e$ in the absence of influence from video or prior live chat messages. By contrast, we denote by $S^{f}(t)$ the time-varying intensity of emotion $f$ in the video, and $\nu^{e,f}$ the cross-influence from emotion $f$ in the video to emotion $e$ in the live chat. The term $\nu^{e,f}\leavevmode\nobreak\ S^{f}(t)$ is thus the time-varying intensity at which emotions of type $f$ in the video triggers emotions of type $e$ in the live chat.
The intensity $S^{f}(t)$ is obtained as an interpolation of observed emotions $f$ in the subtitles of the video. More formally, $S^{f}(t)=\sum_{\tau^{f}_{j}<t}s^{f}_{\tau_{j}}(t)$ where $\left\{\tau^{f}_{j}\right\}$ enumerates all the times at which a subtitle of emotion $f$ appears in the video. Here, $s^{f}_{\tau_{j}}(t)$ captures the time-varying influence of a video subtitle of emotion $f$ appearing in the video at time $\tau_{j}$. Clearly, $s^{f}_{\tau_{j}}(t)=0$ for $t<\tau_{j}$. In order to incorporate the video impact, we need to make reasonable assumptions on the shape of the temporal dependence of the influence of emotions in the video on emotions in the live chat. It is natural to assume that it takes a few seconds for the audience to process the transcript and react. We also expect the influence of the video to peak after a rapid increase in intensity and then fade out with time. Empirical evidence demonstrates that information retention in humans decays with a fat tail, that is, slower than an exponential [67]. To approximately capture the rapid initial increase and subsequent slow decline, we use the log-normal function to parametrize the memory of a subscript in the video:
@@
To fix $\mu$ and $\sigma$, we assume that the intensity of each transcript peaks $2$ seconds after appearance and that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance. In other words, we assume that the maximum of the log-normal function lies at $2$ seconds, $\exp(\mu-\sigma^{2})=2$, and that the median is equal to $10$ seconds, $\exp(\mu)=10$. Numerically solving for $\mu$ and $\sigma$ yields $2.3$ and $1.3$ respectively in units of minutes. To ensure that our results are not strongly dependent on this choice of $\mu$ and $\sigma$, we have checked that our results remain qualitatively similar for different values and functional shapes (SI Appendix). Alternatively, these parameters could be directly fitted from the data, a task which we leave for future research.
In summary, the shape $S^{f}(t)$ is a sum of log-normal functions, with local peaks at $2$ seconds after the appearance of a new subtitle of emotion $f$. An example of $S^{\text{sad}}$ is shown in Figure 1(d). We further stress that our framework is different from the typical multivariate Hawkes process in that expression (2) contains cross-influence terms from different exogenous sources. This is because we condense a bi-multivariate Hawkes model containing two systems of events (video and chat) into one multivariate Hawkes self-excited conditional point process for the chat. This specification is therefore necessary to distinguish video-based emotion events and chat-based emotion events.
Fitting the Parameters of the Point Process
@@
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}^{e}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}(s)}-$
$\quad\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\phi^{e,f}(s)}.$ (4b)
We follow the common assumption that $\phi$ is exponentially decaying and write $\phi^{e,f}(t)=\left.\alpha^{e,f}e^{-t/\gamma^{e}}\right/\gamma^{e}$ with decay time $\gamma^{e}$, such that (4b) can be expanded into
$\log L^{e}\left(\mu^{e}_{0},\nu^{e,f},\alpha^{e,f},\gamma^{e}\right)=$ (5)
$\sum_{i=1}^{N^{e}}\log\left(\mu^{e}_{0}+\sum_{f\in\mathcal{E}}\nu^{e,f}S^{f}(t_{i})+\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\alpha^{e,f}\frac{1}{\gamma^{e}}e^{-\frac{1}{\gamma^{e}}(t_{i}-t_{j}^{f})}\right)$
$-\mu^{e}_{0}T-\sum_{f\in\mathcal{E}}\nu^{e,f}M^{f}_{1}+\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\alpha^{e,f}\left(e^{-\frac{1}{\gamma^{e}}(T-t_{j}^{f})}-1\right),$
@@
We estimate the values for $\mu^{e}_{0}$, $\nu^{e,f}$, $\alpha^{e,f}$, and $\gamma^{e}$ maximizing Equation (5) using Quasi-Newton optimization with constraints. The variable bound constraints are specified as follows: $0\leq\alpha\leq 50$, $10^{-6}\leq\nu\leq 10$, $0.1\leq\gamma\leq 20$, $0\leq\mu\leq 50$.